# Partial identification of principal causal effects under violations of principal ignorability

### v2

@@ 8 Discussion
Appendix A Details of MCMC Sampling
@@
$\begin{split}&\dfrac{1}{(2\pi)^{\frac{3n}{2}}\sigma_{y}^{n}\sigma_{s0}^{n}\sigma_{s1}^{n}(1-\rho^{2})^{\frac{n}{2}}}\exp\bigg{(}-\frac{1}{2}\sum\limits_{i=1}^{n}\bigg{(}(\boldsymbol{\theta}_{y}^{T}\boldsymbol{d}_{i}-y_{i})^{2}/\sigma_{y}^{2}+\\ &\qquad(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})^{T}\boldsymbol{\Sigma}_{s}^{-1}(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})\bigg{)}\bigg{)},\end{split}$ (12)
@@
$\begin{split}&\boldsymbol{\beta}_{t}\sim N(\boldsymbol{\mu}_{\boldsymbol{\beta}_{t}},\Sigma_{\boldsymbol{\beta}_{t}}),\ \lambda_{t}\sim N(\mu_{\lambda_{t}},\sigma_{\lambda_{t}}^{2}),\ \phi_{t}\sim N(\mu_{\phi_{t}},\sigma_{\phi_{t}}^{2}),\ \boldsymbol{\gamma}\sim N(\boldsymbol{\mu}_{\boldsymbol{\gamma}},\Sigma_{\boldsymbol{\gamma}}),\\ &\boldsymbol{\alpha}\sim N(\boldsymbol{\mu}_{\boldsymbol{\alpha}},\Sigma_{\boldsymbol{\alpha}}),\ \sigma_{y}^{2}\sim IG(\eta_{y},\nu_{y}),\ \sigma_{s0}^{2}\sim IG(\eta_{s0},\nu_{s0}),\ \sigma_{s1}^{2}\sim IG(\eta_{s1},\nu_{s1}),\end{split}$
@@
    $\boldsymbol{\theta}_{y}\mid\cdot\sim N\left(\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+\Sigma_{\boldsymbol{\theta}_{y}}^{-1}\bigg{)}^{-1}\bigg{(}\sum\limits_{i=1}^{n}y_{i}\boldsymbol{d}_{i}/\sigma_{y}^{2}+\Sigma_{\boldsymbol{\theta}_{y}}^{-1}\boldsymbol{\mu}_{\boldsymbol{\theta}_{y}}\bigg{)},\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+\Sigma_{\boldsymbol{\theta}_{y}}^{-1}\bigg{)}^{-1}\right),$
@@
    $\boldsymbol{\Sigma}_{\boldsymbol{\theta}_{y}}=\text{diag}(\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{0}},\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{1}},\sigma^{2}_{\lambda_{0}},\sigma^{2}_{\lambda_{1}},\boldsymbol{\Sigma}_{\boldsymbol{\gamma}}).$
@@
    $\boldsymbol{\theta}_{s}\mid\cdot\sim N\left(\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{E}_{i}\boldsymbol{\Sigma}_{s}^{-1}\boldsymbol{E}_{i}^{T}+\Sigma_{\boldsymbol{\theta}_{s}}^{-1}\bigg{)}^{-1}\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{E}_{i}\boldsymbol{\Sigma}_{s}^{-1}\boldsymbol{u}_{i}+\Sigma_{\boldsymbol{\theta}_{s}}^{-1}\boldsymbol{\mu}_{\boldsymbol{\theta}_{s}}\bigg{)},\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{E}_{i}\boldsymbol{\Sigma}_{s}^{-1}\boldsymbol{E}_{i}^{T}+\Sigma_{\boldsymbol{\theta}_{s}}^{-1}\bigg{)}^{-1}\right),$
    where $\boldsymbol{\Sigma}_{\boldsymbol{\theta}_{s}}=\text{diag}(\sigma_{\phi_{0}}^{2},\sigma_{\phi_{1}}^{2},\boldsymbol{\Sigma}_{\boldsymbol{\alpha}})$.
@@
    $\sigma_{y}^{2}\mid\cdot\sim IG\left(\frac{n}{2}+\eta_{y},\sum\limits_{i=1}^{n}(\boldsymbol{\theta}_{y}^{T}\boldsymbol{d}_{i}-y_{i})^{2})/2+\nu_{y}\right).$
@@
      $\begin{split}p(\sigma_{s0}^{2}\mid\cdot)&\propto\sigma_{s0}^{-n-2-2\eta_{s0}}\exp\bigg{(}-\frac{1}{2(1-\rho^{2})\sigma_{s1}^{2}}\sum\limits_{i=1}^{n}\bigg{(}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i0}-s_{i0})^{2}\sigma_{s1}^{2}/\sigma_{s0}^{2}-\\ &\qquad 2\rho\sigma_{s1}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i0}-s_{i0})(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i1}-s_{i1})/\sigma_{s0}\bigg{)}-\frac{\nu_{s0}}{\sigma_{s0}^{2}}\bigg{)}.\end{split}$
@@
      $\begin{split}p(\sigma_{s1}^{2}\mid\cdot)&\propto\sigma_{s1}^{-n-2-2\eta_{s1}}\exp\bigg{(}-\frac{1}{2(1-\rho^{2})\sigma_{s0}^{2}}\sum\limits_{i=1}^{n}\bigg{(}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i1}-s_{i1})^{2}\sigma_{s0}^{2}/\sigma_{s1}^{2}-\\ &\qquad 2\rho\sigma_{s0}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i0}-s_{i0})(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i1}-s_{i1})/\sigma_{s1}\bigg{)}-{\nu_{s1}}/{\sigma_{s1}^{2}}\bigg{)}.\end{split}$
@@
    $\begin{split}\sigma_{s}^{2}\mid IG\bigg{(}n+\eta_{s},\sum\limits_{i=1}^{n}(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})^{T}\boldsymbol{R}_{s}^{-1}(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})/2+\nu_{s}\bigg{)},\end{split}$
@@
    $\begin{split}p(\rho\mid\cdot)&\propto\prod_{i=1}^{n}\bigg{(}(\zeta_{t_{i}}-\psi_{t_{i}}^{2})^{-1/2}\exp\bigg{(}-\frac{1}{2}\sum\limits_{i=1}^{n}\frac{(\psi_{t_{i}}s^{\prime\prime}_{i}/\sigma_{st}-y^{\prime\prime}_{i})^{2}}{\zeta_{t_{i}}-\psi_{t_{i}}^{2}}\bigg{)}\bigg{)},\end{split}$
@@
Appendix B Results for Binary Intermediates
@@
$\begin{split}P(\rho|\mathcal{D})&\propto\prod_{i=1}^{n}(\sum\limits_{s_{i}^{mis}=0,1}\exp(-\frac{(y-\lambda_{t_{i}}-\boldsymbol{\beta}_{t_{i}}^{T}\boldsymbol{u})^{2}}{2\sigma_{y}^{2}})p_{00}^{(1-s_{i0})(1-s_{i1})}p_{01}^{(1-s_{i0})s_{i1}}p_{10}^{s_{i0}(1-s_{i1})}\\ &\qquad p_{11}^{s_{i0}s_{i1}}).\end{split}$
B.1 Identification of the Association Between Principal Strata
@@
Let $f_{t}(y_{i},s_{i})$ denote the conditional marginal pdf of $(Y_{i},S_{i})|\rho,T_{i}=t$, for $t=0,1$. For simplicity, at times we simply refer to this as $f_{t}$. Let $h_{n}(\rho)$ denote the log normalized posterior, defined as $h_{n}(\rho)=\sum\limits_{i=1}^{n}\log(f_{t_{i}}(y_{i},s_{i}|\rho))$, and $q_{n}(\rho)=h_{n}(\rho)/n$. Let $q_{0}(\rho)=P(T_{i}=1)E(\log(f_{1}))+P(T_{i}=0)E(\log(f_{0}))$.
@@
where $\rho=p_{11}$. Similar to Appendix .1.1, we only need to check identification for $q_{0}$ and uniform convergence for $q_{n}$.
@@
    Following in the same spirit as the proof for continuous intermediates in Appendix .1.1, we only need to show that $\text{Var}(\log(f_{T_{i}}(Y_{i},S_{i}|\rho)|T_{i}=t)=\text{Var}(\log f_{t}(Y(0),S(0)))$ and $E(\log(f_{T_{i}}(Y_{i},S_{i}|\rho)|T_{i}=t)=E(\log f_{t}(Y(0),S(0)))$ is bounded w.r.t $\rho$ ($p_{11}$) for $t=0,1$. It is sufficient for us to show that $E\log^{2}f_{t}$ is bounded w.r.t $\rho$ ($p_{11}$) for $t=0,1$.
@@
B.2 Identification up to Sign When Principal Ignorability Fails
@@
Now, all remaining unknown parameters (except for the signs of ($\beta_{01},\beta_{10}$)) can be expressed in terms of $\sigma_{y}$ and the marginal parameters described above. We plug the solutions into conditional marginal pdf of $(Y,S)|T=0$ and obtain $f_{0}(y,s|\sigma_{y}^{2})$ (this can also be done with $f_{1}(y,s|\sigma_{y}^{2})=f_{1}(y,s|(\sigma_{y}^{\ast})^{2})$). Next, we only need to show that $f_{0}(y,s|\sigma_{y}^{2})=f_{0}(y,s|(\sigma_{y}^{\ast})^{2})$, for any $y\in\mathbb{R}$ and $s\in\{0,1\}$, implies that $\sigma_{y}=\sigma_{y}^{\ast}$.
Suppose, for contradiction, that there is $\sigma_{y0}\in\mathbb{R}^{+}$ such that $f_{0}(y,s|\sigma_{y0}^{2})=f_{0}(y,s|(\sigma_{y}^{\ast})^{2})$, for any $y\in\mathbb{R}$ and $s\in\{0,1\}$. Note that
@@
for any $\sigma_{y}\in\mathbb{R}^{+}$. For any $\delta>0$ and $\sigma_{y}\in\{\sigma_{y0},\sigma_{y}^{\ast}\}$, there is $y_{0}\in\mathbb{R}$, such that for $y>y_{0}$, we have
@@
which contradicts the fact that $f_{0}(y,s|\sigma_{y0}^{2})=f_{0}(y,s|(\sigma_{y}^{\ast})^{2})$, for any $y\in\mathbb{R}$ and $s\in\{0,1\}$. This indicates that $\sigma_{y}=\sigma_{y}^{\ast}$ and $\sigma_{y}$ is identifiable. Moreover, $\beta_{01}$ and $\beta_{10}$ are also identifiable (except for the signs). The reason that the signs are not identified or are weakly identified can be seen from the first equation in (10), where knowing $\sigma_{y}$ only identifies $\beta_{t,1-t}^{2}$.
B.3 Simulations with binary intermediate
@@
The trace plots of $\beta_{01}$ and $\beta_{10}$, both with and without the positive sign constraint, are shown in Figures 2 and 3, respectively. Without the sign constraint, the trace plots of $\beta_{01}$ and $\beta_{10}$ tend to explore different signs, and the trace plot of $\beta_{10}$ bounces around the negative of the true value. This highlights our result that these parameters are only identified up to sign, and therefore without a sign constraint, the MCMC algorithm bounces between different values that are equally supported by the data. After applying the positive (true) sign constraint, both trace plots hover around the true values, confirming the identification results for $\beta_{01}$ and $\beta_{10}$ discussed in Appendix B.2.
@@
B.4 Details of MCMC Sampling for binary intermediates
@@
$\begin{split}\dfrac{1}{(2\pi)^{\frac{n}{2}}\sigma_{y}^{n}}\exp\left(-\sum\limits_{i=1}^{n}\frac{(y_{i}-\boldsymbol{\theta}_{y}^{T}\boldsymbol{d}_{i})^{2}}{2\sigma_{y}^{2}}\right)p_{00}^{n_{00}}p_{01}^{n_{01}}p_{10}^{n_{10}}p_{11}^{n_{11}},\end{split}$
where $n_{00}=\sum\limits_{i}(1-s_{i0})(1-s_{i1})$, $n_{01}=\sum\limits_{i}(1-s_{i0})s_{i1}$, $n_{10}=\sum\limits_{i}s_{i0}(1-s_{i1})$, $n_{11}=\sum\limits_{i}s_{i0}s_{i1}$, and
@@
$\boldsymbol{\beta}_{t}\sim N(\boldsymbol{\mu}_{\boldsymbol{\beta}_{t}},\Sigma_{\boldsymbol{\beta}_{t}}),\ \lambda_{t}\sim N(\mu_{\lambda_{t}},\sigma_{\lambda_{t}}^{2}),\ \sigma_{y}^{2}\sim IG(\eta_{y},\nu_{y}),$
@@
    We can update the unobserved values of the intermediate $S(1-t)$ from a Bernoulli with probability
@@
    $\boldsymbol{\theta}_{y}\mid\cdot\sim N\left(\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+\Sigma_{\boldsymbol{\theta}_{y}}^{-1}\bigg{)}^{-1}\bigg{(}\sum\limits_{i=1}^{n}y_{i}\boldsymbol{d}_{i}/\sigma_{y}^{2}+\Sigma_{\boldsymbol{\theta}_{y}}^{-1}\boldsymbol{\mu}_{\boldsymbol{\theta}_{y}}\bigg{)},\bigg{(}\sum\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+\Sigma_{\boldsymbol{\theta}_{y}}^{-1}\bigg{)}^{-1}\right),$
@@
    $\boldsymbol{\Sigma}_{\boldsymbol{\theta}_{y}}=\text{diag}(\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{0}},\boldsymbol{\Sigma}_{\boldsymbol{\beta}_{1}},\sigma^{2}_{\lambda_{0}},\sigma^{2}_{\lambda_{1}}).$
@@
    $\sigma_{y}^{2}\mid\cdot\sim IG\left(\frac{n}{2}+\eta_{y},\sum\limits_{i=1}^{n}(\boldsymbol{\theta}_{y}^{T}\boldsymbol{d}_{i}-y_{i})^{2})/2+\nu_{y}\right).$
@@
    $p(p_{1\cdot},p_{\cdot 1})\propto\prod_{i}\left((1+p_{11}-p_{1\cdot}-p_{\cdot 1})^{(1-s_{i0})(1-s_{i1})}(p_{\cdot 1}-p_{11})^{(1-s_{i0})s_{i1}}(p_{1\cdot}-p_{11})^{s_{i0}(1-s_{i1})}p_{11}^{s_{i0}s_{i1}}\right),$
@@
    $p(p_{11}\mid\cdot)\propto\prod_{i}\left(f_{1}^{t_{i}}(y_{i},s_{i}|p_{11})f_{0}^{1-t_{i}}(y_{i},s_{i}|p_{11})\right).$
@@
Appendix C Additional Plots
We ran additional MCMC chains under the setting described in Section 6.2 with $n$ varying from 300 to 9600. We plotted $\log(n)$ versus the logarithm of the estimated posterior variance, as shown in Fig. 5. This plot shows that when the sample size is large (over 1000), it is nearly linear, indicating that the posterior variance of $\rho$ is approximately $O(n^{-1})$, consistent with the rate stated in Theorem 2.

### v3

@@ A.2.2 Proof of an equivalent expression for Assumption 8
Appendix B Details of MCMC Sampling
@@
[…]
@@
$\begin{split}&\boldsymbol{\beta}_{t}\sim N(\boldsymbol{\mu}_{\boldsymbol{\beta}_{t}},{}_{\boldsymbol{\beta}_{t}}),\ \lambda_{t}\sim N(\mu_{\lambda_{t}},\sigma_{\lambda_{t}}^{2}),\ \phi_{t}\sim N(\mu_{\phi_{t}},\sigma_{\phi_{t}}^{2}),\ \boldsymbol{\gamma}\sim N(\boldsymbol{\mu}_{\boldsymbol{\gamma}},{}_{\boldsymbol{\gamma}}),\\ &\boldsymbol{\alpha}\sim N(\boldsymbol{\mu}_{\boldsymbol{\alpha}},{}_{\boldsymbol{\alpha}}),\ \sigma_{y}^{2}\sim IG(\eta_{y},\nu_{y}),\ \sigma_{s0}^{2}\sim IG(\eta_{s0},\nu_{s0}),\ \sigma_{s1}^{2}\sim IG(\eta_{s1},\nu_{s1}),\end{split}$
@@
    $\boldsymbol{\theta}_{y}\mid\cdot\sim N\left(\bigg(\sumop\displaylimits\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+{}_{\boldsymbol{\theta}_{y}}^{-1}\bigg)^{-1}\bigg(\sumop\displaylimits\limits_{i=1}^{n}y_{i}\boldsymbol{d}_{i}/\sigma_{y}^{2}+{}_{\boldsymbol{\theta}_{y}}^{-1}\boldsymbol{\mu}_{\boldsymbol{\theta}_{y}}\bigg),\bigg(\sumop\displaylimits\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+{}_{\boldsymbol{\theta}_{y}}^{-1}\bigg)^{-1}\right),$
@@
[…]
@@
[…]
@@
    $\sigma_{y}^{2}\mid\cdot\sim IG\left(\frac{n}{2}+\eta_{y},\sumop\displaylimits\limits_{i=1}^{n}(\boldsymbol{\theta}_{y}^{T}\boldsymbol{d}_{i}-y_{i})^{2})/2+\nu_{y}\right).$
@@
      $\begin{split}p(\sigma_{s0}^{2}\mid\cdot)&\propto\sigma_{s0}^{-n-2-2\eta_{s0}}\exp\bigg(-\frac{1}{2(1-\rho^{2})\sigma_{s1}^{2}}\sumop\displaylimits\limits_{i=1}^{n}\bigg((\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i0}-s_{i0})^{2}\sigma_{s1}^{2}/\sigma_{s0}^{2}-\\ &\qquad 2\rho\sigma_{s1}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i0}-s_{i0})(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i1}-s_{i1})/\sigma_{s0}\bigg)-\frac{\nu_{s0}}{\sigma_{s0}^{2}}\bigg).\end{split}$
@@
      $\begin{split}p(\sigma_{s1}^{2}\mid\cdot)&\propto\sigma_{s1}^{-n-2-2\eta_{s1}}\exp\bigg(-\frac{1}{2(1-\rho^{2})\sigma_{s0}^{2}}\sumop\displaylimits\limits_{i=1}^{n}\bigg((\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i1}-s_{i1})^{2}\sigma_{s0}^{2}/\sigma_{s1}^{2}-\\ &\qquad 2\rho\sigma_{s0}(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i0}-s_{i0})(\boldsymbol{\theta}_{s}^{T}\boldsymbol{e}_{i1}-s_{i1})/\sigma_{s1}\bigg)-{\nu_{s1}}/{\sigma_{s1}^{2}}\bigg).\end{split}$
@@
    $\begin{split}\sigma_{s}^{2}\mid IG\bigg(n+\eta_{s},\sumop\displaylimits\limits_{i=1}^{n}(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})^{T}\boldsymbol{R}_{s}^{-1}(\boldsymbol{E}_{i}^{T}\boldsymbol{\theta}_{s}-\boldsymbol{u}_{i})/2+\nu_{s}\bigg),\end{split}$
@@
    $\begin{split}p(\rho\mid\cdot)&\propto\prodop\displaylimits_{i=1}^{n}\bigg((\zeta_{t_{i}}-\psi_{t_{i}}^{2})^{-1/2}\exp\bigg(-\frac{1}{2}\sumop\displaylimits\limits_{i=1}^{n}\frac{(\psi_{t_{i}}s^{\prime\prime}_{i}/\sigma_{st}-y^{\prime\prime}_{i})^{2}}{\zeta_{t_{i}}-\psi_{t_{i}}^{2}}\bigg)\bigg),\end{split}$
@@
Appendix C Results for Binary Intermediates
@@
$\begin{split}P(\rho|\mathcal{D})&\propto\prodop\displaylimits_{i=1}^{n}(\sumop\displaylimits\limits_{s_{i,1-t_{i}}=0,1}\exp(-\frac{(y-\lambda_{t_{i}}-\boldsymbol{\beta}_{t_{i}}^{T}\boldsymbol{u})^{2}}{2\sigma_{y}^{2}})p_{00}^{(1-s_{i0})(1-s_{i1})}p_{01}^{(1-s_{i0})s_{i1}}p_{10}^{s_{i0}(1-s_{i1})}\\ &\qquad p_{11}^{s_{i0}s_{i1}}).\end{split}$
C.1 Identification of the Association Between Principal Strata
@@
Let $f_{t}(y_{i},s_{i})$ denote the conditional marginal pdf of $(Y_{i},S_{i})|\rho,T_{i}=t$, for $t=0,1$. For simplicity, at times we simply refer to this as $f_{t}$. Let $h_{n}(\rho)$ denote the log normalized posterior, defined as $h_{n}(\rho)=\sumop\displaylimits\limits_{i=1}^{n}\log(f_{t_{i}}(y_{i},s_{i}|\rho))$, and $q_{n}(\rho)=h_{n}(\rho)/n$. Let $q_{0}(\rho)=P(T_{i}=1)E(\log(f_{1}))+P(T_{i}=0)E(\log(f_{0}))$.
@@
where $\rho=p_{11}$. Similar to Appendix A.1.1, we only need to check identification for $q_{0}$ and uniform convergence for $q_{n}$.
@@
    Following in the same spirit as the proof for continuous intermediates in Appendix A.1.1, we only need to show that $\text{Var}(\log(f_{T_{i}}(Y_{i},S_{i}|\rho)|T_{i}=t)=\text{Var}(\log f_{t}(Y(0),S(0)))$ and $E(\log(f_{T_{i}}(Y_{i},S_{i}|\rho)|T_{i}=t)=E(\log f_{t}(Y(0),S(0)))$ is bounded w.r.t $\rho$ ($p_{11}$) for $t=0,1$. It is sufficient for us to show that $E\log^{2}f_{t}$ is bounded w.r.t $\rho$ ($p_{11}$) for $t=0,1$.
@@
C.2 Identification up to Sign When Principal Ignorability Fails
@@
[…]
@@
[…]
@@
[…]
C.3 Simulations with binary intermediate
@@
The trace plots of $\beta_{01}$ and $\beta_{10}$, both with and without the positive sign constraint, are shown in Figures 3 and 4, respectively. Without the sign constraint, the trace plots of $\beta_{01}$ and $\beta_{10}$ tend to explore different signs, and the trace plot of $\beta_{10}$ bounces around the negative of the true value. This highlights our result that these parameters are only identified up to sign, and therefore without a sign constraint, the MCMC algorithm bounces between different values that are equally supported by the data. After applying the positive (true) sign constraint, both trace plots hover around the true values, confirming the identification results for $\beta_{01}$ and $\beta_{10}$ discussed in Appendix C.2.
@@
C.4 Details of MCMC Sampling for binary intermediates
@@
$\begin{split}\dfrac{1}{(2\pi)^{\frac{n}{2}}\sigma_{y}^{n}}\exp\left(-\sumop\displaylimits\limits_{i=1}^{n}\frac{(y_{i}-\boldsymbol{\theta}_{y}^{T}\boldsymbol{d}_{i})^{2}}{2\sigma_{y}^{2}}\right)p_{00}^{n_{00}}p_{01}^{n_{01}}p_{10}^{n_{10}}p_{11}^{n_{11}},\end{split}$
where $n_{00}=\sumop\displaylimits\limits_{i}(1-s_{i0})(1-s_{i1})$, $n_{01}=\sumop\displaylimits\limits_{i}(1-s_{i0})s_{i1}$, $n_{10}=\sumop\displaylimits\limits_{i}s_{i0}(1-s_{i1})$, $n_{11}=\sumop\displaylimits\limits_{i}s_{i0}s_{i1}$, and
@@
$\boldsymbol{\beta}_{t}\sim N(\boldsymbol{\mu}_{\boldsymbol{\beta}_{t}},{}_{\boldsymbol{\beta}_{t}}),\ \lambda_{t}\sim N(\mu_{\lambda_{t}},\sigma_{\lambda_{t}}^{2}),\ \sigma_{y}^{2}\sim IG(\eta_{y},\nu_{y}),$
@@
    We can update the unobserved values of the intermediate S(1t)S(1-t) from a Bernoulli with probability
@@
    $\boldsymbol{\theta}_{y}\mid\cdot\sim N\left(\bigg(\sumop\displaylimits\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+{}_{\boldsymbol{\theta}_{y}}^{-1}\bigg)^{-1}\bigg(\sumop\displaylimits\limits_{i=1}^{n}y_{i}\boldsymbol{d}_{i}/\sigma_{y}^{2}+{}_{\boldsymbol{\theta}_{y}}^{-1}\boldsymbol{\mu}_{\boldsymbol{\theta}_{y}}\bigg),\bigg(\sumop\displaylimits\limits_{i=1}^{n}\boldsymbol{d}_{i}\boldsymbol{d}_{i}^{T}/\sigma_{y}^{2}+{}_{\boldsymbol{\theta}_{y}}^{-1}\bigg)^{-1}\right),$
@@
[…]
@@
    $\sigma_{y}^{2}\mid\cdot\sim IG\left(\frac{n}{2}+\eta_{y},\sumop\displaylimits\limits_{i=1}^{n}(\boldsymbol{\theta}_{y}^{T}\boldsymbol{d}_{i}-y_{i})^{2})/2+\nu_{y}\right).$
@@
    $\begin{split}p(p_{1\cdot},p_{\cdot 1})&\propto\prodop\displaylimits_{i}\Big((1+p_{11}-p_{1\cdot}-p_{\cdot 1})^{(1-s_{i0})(1-s_{i1})}(p_{\cdot 1}-p_{11})^{(1-s_{i0})s_{i1}}\\ &\qquad(p_{1\cdot}-p_{11})^{s_{i0}(1-s_{i1})}p_{11}^{s_{i0}s_{i1}}\Big),\end{split}$
@@
    $p(p_{11}\mid\cdot)\propto\prodop\displaylimits_{i}\left(f_{1}^{t_{i}}(y_{i},s_{i}|p_{11})f_{0}^{1-t_{i}}(y_{i},s_{i}|p_{11})\right).$
@@
Appendix D Additional Plots
We ran additional MCMC chains under the setting described in Section 6.1.2 with $n$ varying from 300 to 9600. We plotted $\log(n)$ versus the logarithm of the estimated posterior variance, as shown in Fig. 6. This plot shows that when the sample size is large (over 1000), it is nearly linear, indicating that the posterior variance of $\rho$ is approximately $O(n^{-1})$, consistent with the rate stated in Theorem 2.